# An Entropy-based Text Watermarking Detection Method

### v2

@@ Abstract
Currently, text watermarking algorithms for large language models (LLMs) can embed hidden features to texts generated by LLMs to facilitate subsequent detection, thus alleviating the problem of misuse of LLMs. Although the current text watermarking algorithms perform well in most high-entropy scenarios, its performance in low-entropy scenarios still needs to be improved. In this work, we proposed that the influence of token entropy should be fully considered in the watermark detection process, that is, the weight of each token during watermark detection should be adjusted according to its entropy, rather than setting the weights of all tokens to the same value as in previous methods. Specifically, we proposed an Entropy-based Watermark Detection (EWD) that gives higher-entropy tokens higher weights during watermark detection, so as to better reflect the degree of watermarking. Furthermore, the proposed detection process is training-free and fully automated. In the experiment, we found that our method can achieve better detection performance in low-entropy scenarios, and our method is also general and can be applied to texts with different entropy distributions. Our code and data will be available online.
@@
Yijian LU1, Aiwei Liu2 , Dianzhi Yu1 , Jingjing Li1 , Irwin King1 1The Chinese University of Hong Kong 2Tsinghua University
@@ 1 Introduction
Figure 1: Watermarking algorithm has limited effect on a low-entropy token as compared to a high-entropy token. Therefore, a low-entropy token has a much smaller probability of being watermarked. In a code generation scenario, most of the tokens are low-entropy, resulting in less watermarked tokens and a small detection z-score.
@@
The text watermarking algorithm for LLMs can embed hidden features that can be detected later during the text generation process of LLMs, thus effectively alleviating the problem of misuse of LLMs. The current LLM watermarking technology is usually to embed watermark information in the process of generating tokens during LLM inference. For example, Kirchenbauer et al. (2023a) divides the model’s vocabulary into two lists, green and red, and increases the logit value of tokens in the green list. By modifying the token distributions, it leads the generated text to contain mostly watermarked (green) tokens. In the watermark detection process, the degree of watermark is usually judged by the number of watermarked tokens in the text.
@@
Although current text watermarking algorithms perform well in most high-entropy scenarios, its performance in low-entropy scenarios such as code generation or formatted document generation still needs to be improved. The reason behind this is that low-entropy tokens usually have stronger determinism, are more difficult to modify in the watermarking process, and are easier to affect the quality of the text when forcefully modified Tu et al. (2023).
@@
In figure 1, we illustrate the watermark ratio in tokens of three different entropy levels. A token with higher entropy will have a significantly larger probability to be modified and watermarked. On the contrary, it is rare to see a green token in low-entropy tokens.
@@
When the generated text contains a large number of low-entropy tokens, tokens that can hardly be modified by the watermarking algorithm, the detection results of the watermark detector are more easily affected by the states of these unmodified low-entropy tokens even when the watermark generator managed to modify all high-entropy tokens. For example, the vast low-entropy tokens in a machine-generated text could all appear to be red, and cause the detector to give a low detection score, thus leading to more misjudgments. To address this problem, Lee et al. (2023) proposed a selective watermark detection method named SWEET. Specifically, they only detect tokens that exceed a certain entropy threshold. However, this method has some problems. First, their method requires a manually-set entropy threshold by analyzing samples of code written by humans. At the same time, their method also does not consider the distribution of token entropy, that is, their method treats all tokens that exceed the threshold as the same, while in fact the entropy of these tokens may vastly varies.
@@
To address the aforementioned problems on low-entropy watermark detection without changing the watermark generation process, we introduced an Entropy-based Watermark Detection (EWD) algorithm to accurately reflect the watermark level in the given text. Specifically, we argue that the influence of a token during detection should be proportional to its entropy. For example, the low-entropy tokens should have a limited impact on the detection result. Therefore, we assign an influence weight to each generated token based on its entropy, which is continuous rather than binary as in SWEET. Then, inside the detector, the watermark level of a given text will be the weighted sum of all the watermarked tokens. By creating a gap between the weights of high entropy tokens and low entropy tokens, we ensure that if the generation algorithm failed to watermark a low-entropy token, it will make minimal impact to the detected watermark level. The acquisition of token weight is an independent and training-free process given the whole text entropy, avoiding the trouble of a human dataset or any manual operations. Furthermore, the proposed detection scheme can be applied to machine-generated texts of all entropy levels. There is no need to switch between different detectors facing various kinds of input texts.
@@
  • We proposed an entropy-based watermark detection method called EWD.
@@ 3.2 Text Watermarking
Text watermarking is to embed a piece of message into the given text that can be decoded for the purpose of identification. Here we include a specific branch of it, defined as Watermark for LLMs by Liu et al. (2024), specifically the ones concerning Logits Generation. The general watermarking scheme used in this paper is to modify the logits generation process of an original LLM $M$ to acquire the watermarked LLM, denoted as $\tilde{M}$. The modification on the original logits by adding a small value to the logits of green-list tokens will result in LLM’s preference on those tokens in the output text. Thus such modification could likely produce a watermarked (green) token. The detection algorithm is to calculate the watermark level using the following equation
@@
where $|s|_{G}$ is the number of watermarked tokens and $\gamma$ is the green-list ratio in the model vocabulary. The algorithm would output 1 if the value is significantly greater than a given threshold, and 0 otherwise. In this work, we only concern the watermark detection process.
@@ 3.3 Token Entropy and Low-entropy Scenario
$SE(k)=\sum_{k}{\frac{p_{k}}{1+zp_{k}}},$ (2)
where $p_{k}$ is the probability vector of the token $k$ and $z$ is a scalar. The entropy value achieves minimal if the logits concentrates on a single token and maximal given a uniform distribution.The watermark generation process is more effective on a high-entropy token where the logits distribution is nearly uniform, so that the modification could produce a green-list token easily. On the contrary, a low-entropy token is almost deterministic and hard to be modified. Therefore, the watermarked ratio for a low-entropy token is much lower, resulting in fewer watermarked tokens in the output. A low-entropy scenario refers to the LLM-generated output consisting of a great number of low-entropy tokens. Here we denote the low-entropy token ratio to the whole text as $\varepsilon$.
@@ 4.1 Motivation
In KGW Kirchenbauer et al. (2023a)’s watermark generation method, a positive constant is added to the logits value of all green-list tokens so that the final output would prefer such tokens. The distribution level of the logits, represented by the entropy, matters significantly during this process. When the logits resemble a uniform distribution, meaning that the logits value among tokens are very close, this logits modification would very likely produce a green-list (watermarked) token. An average-high-entropy text would then consist of many watermarked tokens and output a high z-score according to Eq.1. However, when the logits concentrate on one single token, meaning that it is almost deterministic to output that token, this watermark embedding process would produce a watermarked token with a much lower and constant probability. As a result, the number of watermarked tokens and the consequence detection z-score will be much lower for a high-$\varepsilon$ text, for example, the watermarked code in figure 1. This is because to the detector, tokens are of the same influence, no matter how difficult it is to modify them.
@@ 4.2 Proposed Method
As discussed in the previous sections, it is crucial for the detector to consider each token’s entropy. And we took a step further than SWEET: assigning an entropy-based weight to each token, representing their influence in the detection process.
In figure 1, we can observe that a high-entropy token is easier to be modified and thus has a larger probability of being watermarked than a low-entropy token. Since a low-entropy token is hard to modify, then whether it is watermarked or not should have a limited influence on the overall watermark level. Specifically, we argued that when an easy-to-modify token is not watermarked, it should contribute more to showing that the text $T$ is not watermarked than a hard-to-modify token. That is to say, The influence of a non-watermarked token to yielding a negative detection result, $i.e.$, the given text is not watermarked, should be proportional to its likelihood to be modified, and same for the influence of a watermarked token to yielding a positive detection result:
@@
When reflected in the real world, this influence index is represented by the token weight. Token with a larger spike entropy, meaning that it is easier to be modified, will be given a larger influence weight. On the contrary, a low-entropy token will be given a much lower weight close to 0. As the sum of all watermarked token weights, $|s|_{G}$ will receive a smaller numeric loss when observing a non-watermarked low-entropy token than a non-watermarked high-entropy token, and a larger gain when observing a watermarked high-entropy token than a watermarked low-entropy token.
@@
Pseudocode of our proposed detection method is provided in Algorithm 1. We first acquire the text entropy by computing the model logits of each token. The next step is to determine the influence weight of each token, the output of a designed function with their entropy values as inputs. Due to the narrow range of entropy values, our goal here is to create a bigger gap between tokens of high entropy and low entropy. We achieve this by the function $ComputeWeight$, which is training-free and a totally automated process. Here, we first normalize the text entropy. Then we input the processed entropy into a monotonic increasing function, linear for example, to acquire the final weights. After obtaining the entropy-based weights for each token, we then apply the standard detection procedure by Kirchenbauer et al. (2023a) to determine the group of watermarked tokens, $i.e.$, computing the green lists using the detection key and previously generated tokens. Finally, we sum up the weights of the watermarked tokens and calculate the z-statistic using Eq.4 given the sum and weight vector of all detected tokens. The detector would give a watermark-positive result if the z-statistic is greater than the given threshold.
In real-world low-entropy scenarios, With our EWD, the vast majority of non-watermarked low-entropy tokens will not prevent the detector from yielding a positive conclusion. Meanwhile, facing human-written texts, any non-watermarked high-entropy token will strongly suggest the detector to make a negative output.
@@ 5 Theoretical Analysis
To do so, we adopted Kirchenbauer et al. (2023a)’s steps of approach. Each token in both watermarked and non-watermarked texts can either be green or not, so we can model it as a binomial distribution. While Zhao et al. (2023) set window size in the watermark generator as zero and exhibited major token dependence, $i.e.$, the same tokens in the text are of same watermark status (either all green or all red), in our case, where a non-zero window size is adopted, the token dependence is significantly weakened. For example, the same token with different prefix would exhibit different watermark status. Therefore, to simplify and facilitate the process of theoretical analysis, we made the reasonable assumption that the watermark status of each token is independent. Hence, the sum of green tokens in the text can be approximated with a normal distribution $\mathcal{N}(\mu,\sigma^{2})$ with different sets of mean $\mu$ and variance $\sigma^{2}$ for different detection schemes. Finally, we calculate and compare the probability of the sum surpassing or falling behind the detection threshold $|\tilde{s}|_{G}$, which is the $|s|_{G}$ value in each detection function with a fixed $z$.
@@ 5.1 Type-I Error
In a human-written text, each token is independent of the watermark algorithm. Therefore, its probability to be included in the green list is $\gamma$. Therefore, KGW Kirchenbauer et al. (2023a)’s sum of green tokens is approximated by $\mathcal{N}(\gamma T,\gamma(1-\gamma)T)$. For SWEETLee et al. (2023) that only considers high-entropy tokens, such distribution is modelled as $\mathcal{N}(\gamma\tilde{T},\gamma(1-\gamma)\tilde{T})$, where $\tilde{T}$ is the text length excluding low-entropy tokens. Our method EWD added weights to tokens based on entropy, leading to a weighted binomial distribution. We could also approximate it to a normal distribution with mean $\mu=\gamma\sum_{i=m}^{|T|-1}{W}_{i}$ and variance $\sigma^{2}=\gamma(1-\gamma)\sum_{i=m}^{|T|-1}{W_{i}}^{2}$.
@@ 5.2 Type-II Error
Originally, Kirchenbauer et al. (2023a) sampled over 500 watermark-generated news scripts and reported an average spike entropy of 0.807. In this section, we focused specifically on cases with a much lower average spike entropy. Based on our data, we approximated the distribution of spike entropy with a power law distribution $f(x)=ax^{a-1}$, where the parameters are $a=0.106$, $loc=0.566$, $scale=0.426$. Under this distribution, the average spike entropy is 0.607. The following probability from Kirchenbauer et al. (2023a) showed a lower bound of the probability of a token $k$ being watermarked by a watermark generator:
@@ 6.1 Experiment Settings
Figure 2: The left figure shows the z-scores of watermarked and human texts in the Rotowire, HumanEval and MBPP datasets respectively, each being detected with 3 different methods. The right figure shows the relationship between token weights and the probability of being watermarked in both watermarked and human texts.
Figure 3: We utilized a variety of weight functions and measured their performance in the code detection datasets. The left figure is a visualization of all studied functions, with normalized spike entropy as input. The right figure shows each function’s detection F1 under 5% FPR with comparison to the SWEET baseline. Each data point can correspond to the illustrated function on the left figure by looking at its shape and x-axis value.

### v3

@@ Abstract
Currently, text watermarking algorithms for large language models (LLMs) can embed hidden features to texts generated by LLMs to facilitate subsequent detection, thus alleviating the problem of misuse of LLMs. Although the current text watermarking algorithms perform well in most high-entropy scenarios, its performance in low-entropy scenarios still needs to be improved. In this work, we proposed that the influence of token entropy should be fully considered in the watermark detection process, that is, the weight of each token during watermark detection should be customized according to its entropy, rather than setting the weights of all tokens to the same value as in previous methods. Specifically, we proposed an Entropy-based Text Watermark Detection (EWD) that gives higher-entropy tokens higher influence weights during watermark detection, so as to better reflect the degree of watermarking. Furthermore, the proposed detection process is training-free and fully automated. In the experiment, we found that our method can achieve better detection performance in low-entropy scenarios, and our method is also general and can be applied to texts with different entropy distributions. Our code and data is available on https://github.com/luyijian3/EWD. Additionally, our algorithm could be accessed through MarkLLMPan et al. (2024).
@@
Yijian LU1, Aiwei Liu2 , Dianzhi Yu1 , Jingjing Li1 , Irwin King1 1The Chinese University of Hong Kong 2Tsinghua University luyijian@link.cuhk.edu.hk, liuaw20@mails.tsinghua.edu.cn king@cse.cuhk.edu.hk
@@ 1 Introduction
Figure 1: Watermarking algorithm has limited effect on a low-entropy token as compared to a high-entropy token. Therefore, it is almost random for a low-entropy token to be green. In a code generation scenario, most of the tokens are low-entropy, resulting in less green tokens and a small detection z-score.
@@
The text watermarking algorithm for LLMs can embed hidden features that can be detected later during the text generation process of LLMs, thus effectively alleviating the problem of misuse of LLMs. The current LLM watermarking technology is usually to embed watermark information in the process of generating tokens during LLM inference. For example, KGW Kirchenbauer et al. (2023a) divides the model’s vocabulary into two lists, green and red, and increases the logit value of tokens in the green list. By modifying the token distributions, it imposes a preference on the green list to the model and leads the model to sample most tokens from the green list. In the detection process, the degree of watermark is usually judged by the number of green tokens in the text.
@@
Although current text watermarking algorithms perform well in most high-entropy scenarios, its performance in low-entropy scenarios such as code generation or formatted document generation still needs to be improved. The reason behind this is that low-entropy tokens usually have stronger determinism, and the modified logits can hardly affect the probability of sampling the tokens from the green list. At the same time, if logits is forcefully modified, the quality of the watermarked text is easier to degrade Tu et al. (2023).
@@
In figure 1, we illustrate the green-list ratio in tokens of three different entropy levels. A token with higher entropy will have a significantly larger probability to be sampled from a green list. On the contrary, it is almost random to see a green token in low-entropy tokens if the green list size is half of the model vocabulary size.
@@
When the generated text contains a large number of low-entropy tokens, tokens that can hardly be affected by the modified logits, the detection results of the watermark detector are more easily affected by the states (colors) of these low-entropy tokens even when the watermark generator managed to sample all high-entropy tokens from the green list. For example, the vast low-entropy tokens in a machine-generated text could all appear to be red, and cause the detector to give a low detection score, thus leading to more misjudgments. To address this problem, Lee et al. (2023) proposed a selective watermark detection method named SWEET. Specifically, they only detect tokens that exceed a certain entropy threshold. However, this method has some problems. First, their method requires a manually-set entropy threshold by analyzing samples of code written by humans. At the same time, their method also does not consider the distribution of token entropy, that is, their method treats all tokens that exceed the threshold as the same, while in fact the entropy of these tokens may vastly varies.
@@
To address the aforementioned problems on low-entropy watermark detection without changing the watermark generation process, we introduced an Entropy-based Text Watermark Detection (EWD) algorithm to accurately reflect the watermark level in the given text. Specifically, we argue that the influence of a token during detection should be proportional to its entropy. For example, the low-entropy tokens should have a limited impact on the detection result. Furthermore, in order to fully reflect this positive relation between token entropy and detection influence, we utilized a monotonically-increasing and continuous function to generate influence weights from token entropy. Then, inside the detector, the watermark level of a given text will be the weighted sum of all the green tokens. By creating a gap between the weights of high entropy tokens and low entropy tokens, we ensure that if the generation algorithm can hardly affect a low-entropy token, the state of that token will make minimal impact to the detected watermark result. The acquisition of token weight is an independent and training-free process given the whole text entropy, avoiding the trouble of a human dataset or any manual operations. Furthermore, the proposed detection scheme can be applied to machine-generated texts of all entropy levels. There is no need to switch between different detectors facing various kinds of input texts.
@@
  • We proposed an entropy-based text watermark detection method called EWD.
@@ 3.2 Text Watermarking
Text watermarking is to embed a piece of message into the given text that can be decoded for the purpose of identification. Here we include a specific branch of it, defined as Watermark for LLMs by Liu et al. (2024), specifically the ones concerning Logits Generation. The general watermarking scheme used in this paper is called KGW Kirchenbauer et al. (2023a), which is to modify the logits generation process of an original LLM $M$ to acquire the watermarked LLM, denoted as $\tilde{M}$. The modification on the original logits by adding a small value to the logits of green-list tokens will result in LLM’s preference on those tokens in the output text. Thus such modification could likely produce a green token. The detection algorithm is to calculate the watermark level using the following equation
@@
where $|s|_{G}$ is the number of green tokens and $\gamma$ is the green-list ratio in the model vocabulary. The algorithm would output 1 if the value is significantly greater than a given threshold, and 0 otherwise. In this work, we only concern the watermark detection process.
@@ 3.3 Token Entropy and Low-entropy Scenario
$SE(k)=\sum_{v}{\frac{p_{v}}{1+zp_{v}}},$ (2)
where $p$ is the logits when sampling the token $k$ and $z$ is a scalar. The entropy value achieves minimal if the logits concentrates on a single token and maximal given a uniform distribution.The watermark generation process is more effective on a high-entropy token where the logits distribution is nearly uniform, so that the modification could produce a green-list token easily. On the contrary, a low-entropy token is almost deterministic and hard to be modified. Therefore, the green-list ratio for a low-entropy token is much lower, resulting in fewer green tokens in the output. A low-entropy scenario refers to the LLM-generated output consisting of a great number of low-entropy tokens. Here we denote the low-entropy token ratio to the whole text as $\varepsilon$.
@@ 4.1 Motivation
In KGW Kirchenbauer et al. (2023a)’s watermark generation method, a positive constant is added to the logits value of all green-list tokens so that the final output would prefer such tokens. The distribution level of the logits, represented by the entropy, matters significantly during this process. When the logits resemble a uniform distribution, meaning that the logits value among tokens are very close, this logits modification would very likely produce a green-list token. An average-high-entropy text would then consist of many green tokens and output a high z-score according to Eq.1. However, when the logits concentrate on one single token, meaning that it is almost deterministic to output that token, this watermark embedding process would hardly affect the token sampling and produce a green token with a much lower probability close to the green list ratio $\gamma$. As a result, the number of green tokens and the consequence detection z-score will be much lower for a high-$\varepsilon$ text, for example, the watermarked code in figure 1. This is because to the detector, tokens are of the same influence, no matter how difficult it is for the logits modification to affect them.
@@ 4.2 Proposed Method
In figure 1, we can observe that a high-entropy token is easier to be affected by the modified logits, and has a larger probability of being sampled from the green list. Since a low-entropy token is hard to affect, then whether it is sampled from the green list or not should have a limited influence on the overall watermark level. Specifically, we argued that when an easy-to-affect (high-entropy) token is not green, it should contribute more to showing that the text $T$ is not watermarked than a hard-to-affect (low-entropy) token. That is to say, The influence of a red token to yielding a negative detection result, $i.e.$, the given text is not watermarked, should be proportional to its entropy, and same for the influence of a green token to yielding a positive detection result:
@@
When reflected in the real world, this influence index is represented by the token weight. Token with a larger spike entropy, meaning that it is easier to be affected, will be given a larger influence weight. On the contrary, a low-entropy token will be given a much lower weight close to 0. As the sum of all green token weights, $|s|_{G}$ will receive a smaller numeric loss when observing a red low-entropy token than a red high-entropy token, and a larger gain when observing a green high-entropy token than a green low-entropy token.
@@
Now we are aware of the positive influence-entropy relation as suggest Eq.3. We decided to take a step further than SWEET: customizing influence weight according to token entropy to fully reflect the influence-entropy relation. While in SWEET, the weight of 1 for above-threshold tokens and the weight of 0 for below-threshold tokens partially reflect the relation, there is not reflection inside the entropy range of above-threshold tokens. Therefore, we utilized a monotonically-increasing and continuous function $f$ to generate weights from token entropy. The weight assigning function can be illustrated as follows:
$W(t)=f(SE(t)-C_{0}).$ (5)
Pseudocode of our proposed detection method is provided in Algorithm 1. We first acquire the text entropy by computing the model logits of each token. The next step is to determine the influence weight of each token, the output of a designed function with their entropy values as inputs. Due to the narrow range of entropy values, our goal here is to create a bigger gap between tokens of high entropy and low entropy. We achieve this by the function $ComputeWeight$, which is training-free and a totally automated process. Here, we first normalize the text entropy. Then we input the processed entropy into a monotonic increasing function, linear for example, to acquire the final weights. After obtaining the entropy-based weights for each token, we then apply the standard detection procedure by Kirchenbauer et al. (2023a) to determine the group of green tokens, $i.e.$, computing the green lists using the detection key and previously generated tokens. Finally, we sum up the weights of the green tokens and calculate the z-statistic using Eq.4 given $|s|_{G}$ and weights of all detected tokens. The detector would give a watermark-positive result if the z-statistic is greater than the given threshold.
In real-world low-entropy scenarios, With our EWD, the vast majority of low-entropy red tokens will not prevent the detector from yielding a positive conclusion. Meanwhile, facing human-written texts, any high-entropy red token will strongly suggest the detector to make a negative output.
@@ 5 Theoretical Analysis
To do so, we adopted Kirchenbauer et al. (2023a)’s steps of approach. Each token in both watermarked and non-watermarked texts can either be green or not, so we can model it as a binomial distribution. While Zhao et al. (2023) set window size in the watermark generator as zero (one universal green list for all tokens) and exhibited major token dependence, $i.e.$, the same tokens in the text are of the same color (either all sampled from the green list or all from the red list), in our case, where a non-zero window size is adopted, the token dependence is significantly weakened. For example, the same token with different prefix would exhibit different sampling result. Therefore, to simplify and facilitate the process of theoretical analysis, we made the reasonable assumption that the sampling of each token is independent. Hence, the sum of green tokens in the text can be approximated with a normal distribution $\mathcal{N}(\mu,\sigma^{2})$ with different sets of mean $\mu$ and variance $\sigma^{2}$ for different detection schemes. Finally, we calculate and compare the probability of the sum surpassing or falling behind the detection threshold $|\tilde{s}|_{G}$, which is the $|s|_{G}$ value in each detection function with a fixed $z$.
@@ 5.1 Type-I Error
In a human-written text, each token is independent of the watermark algorithm. Therefore, its probability to be included in the green list is $\gamma$. Therefore, KGW Kirchenbauer et al. (2023a)’s sum of green tokens is approximated by $\mathcal{N}(\gamma T,\gamma(1-\gamma)T)$. For SWEETLee et al. (2023) that only considers high-entropy tokens, such distribution is modelled as $\mathcal{N}(\gamma\tilde{T},\gamma(1-\gamma)\tilde{T})$, where $\tilde{T}$ is the text length excluding low-entropy tokens. Our method EWD assigned weights to tokens based on entropy, leading to a weighted binomial distribution. We could also approximate it to a normal distribution with mean $\mu=\gamma\sum_{i=m}^{|T|-1}{W}_{i}$ and variance $\sigma^{2}=\gamma(1-\gamma)\sum_{i=m}^{|T|-1}{W_{i}}^{2}$.
@@ 5.2 Type-II Error
Originally, Kirchenbauer et al. (2023a) sampled over 500 watermark-generated news scripts and reported an average spike entropy of 0.807. In this section, we focused specifically on cases with a much lower average spike entropy. Based on our data, we approximated the distribution of spike entropy in low-entropy watermarked texts with a power law distribution $f(x)=ax^{a-1}$, where the parameters are $a=0.106$, $loc=0.566$, $scale=0.426$. Under this distribution, the average spike entropy is 0.607. The following probability from Kirchenbauer et al. (2023a) showed a lower bound of the probability of a token $k$ being sampled from the green list:
@@ 6.1 Experiment Settings
Figure 2: The left figure shows the z-scores of watermarked and human texts in the Rotowire, HumanEval and MBPP datasets respectively, each being detected with 3 different methods. The right figure shows the relationship between token weights and the probability of being sampled from the green list in both watermarked and human texts.
Figure 3: We utilized a variety of weight functions and measured their performance in the code detection datasets. The left figure is a visualization of all studied functions, with normalized spike entropy as input. The right figure shows each function’s detection F1 under 1% FPR with comparison to the SWEET baseline. Each data point can correspond to the illustrated function on the left figure by looking at its shape and x-axis value.